OpenStudy (anonymous):

is X+1 a factor of 3x^5 + 4x^2 - 7?

OpenStudy (anonymous):

If x+1 is a factor, then x=-1 will equate to 0. 3(-1)^5 + 4(-1)^2 - 7 = 3(-1) + 4(1) - 7 = -3 + 4 - 7 = -6. So, no, it is not a factor.

OpenStudy (anonymous):

Thank u so very much! God bless